The length of a rectangle is 4 cm more than the width. The area is 45 cm ^ 2. Find the length and width.
Answer by venugopalramana(3286)

The length =L SAY of a rectangle is 4 cm more than the width =W SAY. The area is 45 cm ^ 2. Find the length and width.
L=W+4
AREA = LW = (W+4)W =45
W^2+4W-45=0
W^2+9W-5W-45=0
W(2+9)-5(W+9)=0
(W-5)(W+9)=0
W-5=0...OR...W=5
HENCE L=4+5=9